OpenStudy (anonymous):

please help will give medal! need help distributing f(x)=(x-3)(x-2) and h(x)= (x-2i)(x+2i)

OpenStudy (anonymous):

\[(a+b)(c+d) = ac+ad+bc+bd\] Setting b=ki and d=li gives: \[(a+mi)(c+di) = ac + a*li + ki*c + ki*li = ac - kl + (al + kc)i\] Setting c = a and l = -k gives: \[(a+ki)(a-ki) = a^2 + k^2 + (-ak + ka)i = a^2 + k^2\] This is a fundamental result.